no

How many rows and columns will the above normal form payoff matrix representation have?
Answer #1

Player 1 moves at 2 information set ,there are 2 actions at each set .

So, the number of strategies of P1= 2*2= 4

Player 2 moves at 1 information set , there are 2 actions at each set.

So, the number of strategies of P2= 1*2= 2

Number of rows equals number of P1 strategies and number of column equals the number of P2 strategies.

Number of rows=4

Number of column= 2